OpenStudy (anonymous):

Sammy ate ten less than three times as many strawberries as Amy ate. Sammy ate 14 strawberries. Choose the equation that you would use to find the number, n, of strawberries that Amy ate.

OpenStudy (anonymous):

A. 14 = 3(n - 10) B. 10 = 3(n - 14) C. 10 = 3n - 14 D. 14 = 3n - 10

OpenStudy (jadeishere):

3n - 10 = 14

OpenStudy (jadeishere):

(the number of berries Amy ate was 8, by the way >.<

OpenStudy (anonymous):

so it is a

OpenStudy (jadeishere):

D

OpenStudy (anonymous):

preciate it